# One VLM to Keep it Learning: Generation and Balancing for Data-free Continual Visual Question Answering

### v1

@@ Abstract
Vision-Language Models (VLMs) have shown significant promise in Visual Question Answering (VQA) tasks by leveraging web-scale multimodal datasets. However, these models often struggle with continual learning due to catastrophic forgetting when adapting to new tasks. As an effective remedy to mitigate catastrophic forgetting, rehearsal strategy uses the data of past tasks upon learning new task. However, such strategy incurs the need of storing past data, which might not be feasible due to hardware constraints or privacy concerns. In this work, we propose the first data-free method that leverages the language generation capability of a VLM, instead of relying on external models, to produce pseudo-rehearsal data for addressing continual VQA. Our proposal, named as GaB, generates pseudo-rehearsal data by posing previous task questions on new task data. Yet, despite being effective, the distribution of generated questions skews towards the most frequently posed questions due to the limited and task-specific training data. To mitigate this issue, we introduce a pseudo-rehearsal balancing module that aligns the generated data towards the ground-truth data distribution using either the question meta-statistics or an unsupervised clustering method. We evaluate our proposed method on two recent benchmarks, i.e. VQACL-VQAv2 and CLOVE-function benchmarks. GaB outperforms all the data-free baselines with substantial improvement in maintaining VQA performance across evolving tasks, while being on-par with methods with access to the past data. Code will be made public soon.
@@ 1 Introduction
We evaluate GaB on continual learning VQA benchmarks, namely VQACL-VQAv2 [59] and CLOVE-function [22] and compare the proposed approach with respect to state-of-the-art and continual learning baselines. Results show a large margin improvement on all considered settings among data-free methods. We also provide thorough analysis for each proposed element and verify our proposed method can achieve real-data alike distribution in the Q&A generation, being effective with less memory buffer.
@@ 4.1 Comparison
GaB-classifier 47.65 3.61
GaB-clustering 48.40 1.40
@@ 4.2 Ablation studies
GaB generates questions freely, based on the image content. As we discussed in Sec. 3.2, this results in a skewed distribution towards more prevalent question types and we need to fix this issue via a post-generation balancing strategy. A different direction is to avoid this problem altogether by conditioning the question generation process on target question types. We study this alternative strategy, and compare the proposed approach with GaB-conditioning where the language model is trained by explicitly prompting to generate questions based on their type, p=”Question type <type>: Question: <question> Answer: <answer>”. During inference, question generation is similarly driven by the classifier partitioning function to ensure a balanced distribution. Table 3 presents the results of this conditioning approach in comparison to our methods involving only post-generation processing. We observe that the results of GaB-conditioning are inferior to ours, going from 47.65 to 44.07 in terms of AP. Upon further inspection, we find that forcing the model to produce questions related to certain types can lead to false assumptions, such as forcing to ask a question type that is not relevant for the given image (see Fig. 5 for illustration). As a consequence, the constructed rehearsal buffer can be noisy, thus potentially hampering the learning process.
@@
In our approach, the rehearsal generation simultaneously produces question-answer pairs based on the image input. To explore alternative strategies, we investigate the impact of using our trained answer generation model to respond to these generated questions in a self-training fashion (GaB-self). This experiment aims to assess whether there is a benefit to decoupling the generation of questions and answers during the rehearsal phase. The results of this study are detailed in Table 3. As can be observed, GaB-self gets a noticeable  -2.65 drop in AP by using pseudo-answers from the answer generation module instead of generating answers concurrently with the questions via the QA language projection head. We hypothesize that this performance drop is due to the quality of the answers provided by the answer generation model. Specifically, the answers tend to be noisy or incomplete, inducing to confirmation bias as the model is fine-tuned on these inaccurately labeled answers. Some qualitative examples of GaB-self can be observed in the bottom row of Fig. 5.
@@
In this study, we aim to test the impact of generating question-answer (QA) pairs using images from the current task versus images from previous tasks. The motivation for this analysis is to understand why our rehearsal strategy, which leverages current task images for QA pair generation, consistently outperforms traditional real-data rehearsal methods. We verify our finding in Table 3 where GaB-pastimages applies the same question-answer generation strategy to past visual images. The results highlight the performance difference of around 10% compared to the rehearsal using current task images, as in GaB-classifier. We hypothesize that generating QA pairs on current task images provides richer contextual relevance for the model, akin to augmentation. As multiple questions are asked on the same current images, more context is provided aiding the model to accurately answer questions. Moreover, by engaging with current images, the model bridges the gap between the different tasks encountered, providing an effective regularization for positive transfer. Hence, the model reinforces and integrates new knowledge with previously learned information in a more meaningful way, enhancing its ability to retain relevant skills across tasks.
@@ 5 Conclusion
Appendix A Supplementary Overview
In this supplementary material we present further results and ablation analysis on the presented method GaB, implementation details and qualitative visualization of generated question-answer pairs. Specifically, while in Section B, we report extensive results on the method, Section C ablates the proposed pseudo-rehearsal strategy. We continue with Section D and Section E describing in further detail the datasets and the implementation details of the approach, respectively. Finally, Section F concludes by providing qualitatives of generated question answer pairs.
Appendix B Extended Results
We here report extended empirical results for the proposed method GaB, including the intermediate evaluation of the sequentially training model and the robustness of the approach to different task orders.
B.1 Per-Task Performance Analysis
We evaluate the performance of the learning VQA model on the different tasks as training progresses.
Figure 6: Per-task performance in terms of AP across different tasks in the VQACL-VQAv2 benchmark.
Figure 7: Per-task performance in terms of AP across different tasks in the CLOVE-function benchmark.
Figure 6 showcases the performance variations across various tasks within the VQACL-VQAv2 benchmark in terms of AP, highlighting how each continual learning method adapts over time. As can be noted, despite building on pseudo-rehearsal samples only, throughout the entire sequential adaptation GaB achieves accuracy on par with the rehearsal strategy that leverage past real data. A similar behaviour is observed in Figure 7 detailing the sequential task performance for the CLOVE-function benchmark during sequential training.
B.2 Continual Learning Across Task Orders on CLOVE
Continual learning performance can significantly vary depending on the order in which tasks are presented. To explore this variability, we evaluate GaB across three different task orders of the CLOVE-function benchmark. Let us denote each task with its initial letter, i.e., Objects (o), Attributes (a), Relations (r), Logical (l) and Knowledge (k), we consider task orders ‘oarlk’, ‘rolak’, and ‘lkora’.
Table 4: The continual learning performance in terms of AP and AF on the CLOVE-function dataset considering different task orders.
oarlk lkora rolak
Method AP ($\uparrow$) AF ($\downarrow$) AP ($\uparrow$) AF ($\downarrow$) AP ($\uparrow$) AF ($\downarrow$)
Multi-task 32.26
[…]
Seq-FT 22.70 22.19 13.35 24.24 12.35 23.61
GaB w/o balancing 37.01 3.61 25.48 16.9 29.18 7.97
GaB-clustering (Ours) 40.70 1.40 31.59 10.64 26.17 9.67
The results are summarized in Table 4 in terms of Average Performance (AP) and Average Forgetting (AF). Despite the differences in AP and AF across the task sequences, similar behaviours are observed: our approach GaB-clustering performs competitively with the rehearsal strategy and it provides a large margin improvement to the Seq-FT baseline. Notably, the ‘oarlk’ sequence consistently shows better performance metrics compared to ‘lkora’ and ‘rolak’, suggesting that the order in which tasks are encountered can influence the efficacy of the learning process. In the ‘oarlk’ sequence, our GaB-clustering method demonstrated the best resilience against forgetting with an AF of 1.40, and a high AP of 40.70, underscoring its robustness in handling the challenges presented by this particular sequence. We see a similar trend in the sequence ‘lkora’ where our method has the best overall AP and AF. Conversely, in the ‘rolak’, GaB w/o balancing shows the most effective, indicating that pseudo-generation aids forgetting mitigation while the clustering-based balancing strategy might require more careful hyperparameter tuning, such as the number of clusters to use.
B.3 Balancing Questions
We visualize the question types distribution of rehearsal samples before (Generated) and after balancing (Balanced) with our pseudo-rehearsal balancing module, compared to the ground truth one (Real). These visualizations help illustrate the impact of our balancing technique on the diversity of question types generated during the pseudo-rehearsal data generation.
Figure 8: Question distribution before and after pseudo-rehearsal balancing for the VQACL-VQAv2 benchmark. The figure shows the distribution across different question categories for old tasks (as per plot title) generated from last task count visual images.
Figure 9: Question distribution before and after pseudo-rehearsal balancing for the CLOVE-function benchmark. The figure shows the distribution across different question categories for old tasks (as per plot title) generated from last task knowledge visual images.
Figure 8 illustrates the distribution alignment for VQACL-VQAv2 demonstrating how GaB-clustering ensures no single question type dominates the training process. We show similar results on CLOVE-function in Figure 9.
Appendix C Further ablations
We further ablate GaB on the number of balancing clusters and the use of the question-answer generation module for dynamic sampling of the rehearsal data.
Figure 10: The continual learning performance of GaB-clustering in terms of AP when varying the number of balancing clusters $K$ on the VQACL-VQAv2 dataset.
C.1 Varying the number of balancing clusters
We explore the impact of varying the number of clusters in our balanced cluster strategy GaB-clustering. We ablate to determine the optimal number of clusters that yields the best performance. The results of this study are illustrated in Figure 10, which displays the average precision achieved across different cluster counts $K$. Our findings indicate that setting the number of clusters to 7 maximizes average perfomance.
C.2 Dynamic Pseudo-Rehearsal
Traditional rehearsal strategies for continual learning are constrained in the buffer dimension due to the limited availability of question-answers from previous tasks. Thanks to the dynamic generation of samples, pseudo-rehearsal strategies could possibly rely on larger replay buffers instead. We here explore a dynamic version of the presented approach (GaB-dynamic) where pseudo-samples are generated on-the-fly based on current task data. At each training step this method involves dynamically generating a question related to previous tasks for each current batch visual image. As the alignment of the question distributions is non-trivial in the batch low-samples setting, no balancement to the generated questions-answer pairs is applied.
Table 5 compares the approaches in terms of AP and AF for the CLOVE-function setting. Despite the larger number of replay samples available to GaB-dynamic, the results indicate that GaB-clustering achieves the highest performance with a +4.53% and -3% in the AP and AF metrics, respectively. Intuitively, we observe limited variability of QA samples that could be generated within each batch with a consequent drop on the final performance.
Table 5: The continual learning performance in terms of AP and AF on the CLOVE-function dataset.
Method AP ($\uparrow$) AF ($\downarrow$)
GaB-dynamic 36.17 5.26
GaB-classifier (Ours) 37.97 5.25
GaB-clustering (Ours) 40.70 2.26
Appendix D Dataset Details
We provide comprehensive details on both datasets utilized in our study. The VQACL-VQAv2 benchmark is comprised of 10 distinct tasks, each with its own set of questions tailored to specific aspects of visual and textual understanding. The different tasks considered are in order: Recognition, Location, Judge, Commonsense, Count, Action, Color, Type, Subcategory and Causal. We refer the reader to the original paper for extensive details on the benchmark. On the other hand, the CLOVE benchmark includes 6 tasks. For fair comparison, we restrict our evaluation to only 5 tasks due to the specialized architecture and auxiliary features needed for answering scene-text questions asking to OCR present text.
Each task within these datasets is associated with questions that are categorized into various types depending on their content and focus. These question types are labeled systematically to facilitate targeted training and analysis. While in VQACL-VQAv2 meta-information on question types models the initial words used in the question construction, e.g. ”what type”, ”is the”, ”where is”, ”how many”, differently, CLOVE-function auxiliary information models the property being queried and the expected answer, for instance ”MaterialChoose”, ”activityWho” or ”relVerify”.
Appendix E Implementation details
We implement our strategy in PyTorch [31] and employ the Hugging Face [^1] https://huggingface.co/ implementation of the BLIP-2[23] architecture, specifically the opt-2.7b version. For textual generation, we follow standard practice and fix the max_new_tokens for both generated answers (2 tokens) and pseudo QA pairs (20 tokens). The repetition penalty is set to 1.2. In line with prior work, we prompt BLIP-2 for answer generation with the prompt p="Question: <question> Answer:", while pseudo-QA has an empty prompt p="", akin to the original BLIP-2 strategy.
Baselines.
We implemented continual learning strategies following an open-source codebase for CL[27]. The strategies are applied to the same BLIP-2 architecture as GaB. For the regularization-based methods, Elastic Weight Consolidation (EWC) and Memory Aware Synapses (MAS), we set the regularization parameters to 1.0 and compute importance weights on the token classifier generating the output textual sequence. For the Learning to Prompt (L2P) approach the prompting strategy is applied to the visual encoder only. We adopt different configurations depending on the benchmark: for VQACL-VQAv2, we utilize a prompt pool size of 10 with a regularization parameter of 1.0, whereas for CLOVE-function, the prompt pool size is increased to 50 and the regularization parameter is adjusted to 0.5. These settings are carefully chosen to optimize performance across the diverse conditions presented by each benchmark.
Appendix F Qualitatives Results
We report qualitative results on the generated question-answer pairs providing both positive and negative examples. In Figure 11-a and Figure 11-b GaB successfully generates both accurate and contextually appropriate questions and accompanying answers. For instance, questions such as “What material is the floor made of?” with the answer “tile” and “Is there a bear in the picture?” with the answer “yes” demonstrate the model’s ability to understand and respond correctly based on the visual data. Differently, in Figure 11-c we highlight a scenario where the generated question is vague and lacks specificity. An example from this row includes a question like “What color is the jacket?” which, although correct (answered as “orange”), does not specify which jacket to look for as there are multiple people wearing jackets in the image. Finally, Figure 11-d presents instances where the model’s generated answers are incorrect. For example, the question “How many planes are there?” receives the answer “3”, whereas we can clearly see there are in fact only 2 airplanes, indicating the model’s challenges in some contexts or its misinterpretation of the visual content.
Figure 11: Qualitatative results of GaB generated QA pairs on VQACL-VQAv2 datasets. (a)-(b) Correctly generated questions answer pairs; (c) question-answer pairs with ill-posed questions; (d) question-answer pairs with wrong generated answers.
Appendix G Complexity overhead discussion
Table 6: Computational overhead in terms of training parameters and time. Param count indicates the fraction of trained parameters, TFlops indicates Tera Flops in forward+backward passes in 1 epoch, Computational time metrics report the processing time requirements (in seconds) per one epoch of a single task in CLOVE-function across different methods.
Method Param count TFlops Computational time (s)
Training Generation Balancing
Rehearsal 0.05 2711.05 35,085 n.a. n.a.
LAMOL* 0.1 5670.19 49,650 363 n.a.
GaB-clustering (Ours) 0.3 3568.38 42,885 1,452 273
As a data-free rehearsal strategy  GaB trades-off the prohibited access to past tasks data with increased computational complexity for sample generation. More precisely, GaB requires two passes of the employed VLM: a first forward-backward pass trains the qa projection layer $(f_{v\to qa})$ and answer projection layer $(f_{v\to a})$ while the second forward pass allows to generate rehearsal data pairs. Limited overhead is required for balancing the questions according to precomputed real-data clustering statistics, with no need for further modification. Only the task-specific projection heads and answering head are trained, while the rest of the model remains frozen, making the process computationally efficient. We evaluate the computational overhead for pseudo-rehearsal generation and balancing of the presented approach, results are presented in Tab. 6, where we can see that the approach falls behind the rehearsal baseline while being competitive with the data-free version of LAMOL*. Both GaB and LAMOL* suffer from the longer training time due to the requirements of learning how to generate question-answer pairs, however, GaB avoids replaying for the generation task, ending up with shorter training time at the cost of larger number of training parameters. No additional computation is required at inference time where task-specific heads can be discarded and the single shared answering head solves for the VQA task.

### v2

@@ Abstract
Vision-Language Models (VLMs) have shown significant promise in Visual Question Answering (VQA) tasks by leveraging web-scale multimodal datasets. However, these models often struggle with continual learning due to catastrophic forgetting when adapting to new tasks. As an effective remedy to mitigate catastrophic forgetting, rehearsal strategy uses the data of past tasks upon learning new task. However, such strategy incurs the need of storing past data, which might not be feasible due to hardware constraints or privacy concerns. In this work, we propose the first data-free method that leverages the language generation capability of a VLM, instead of relying on external models, to produce pseudo-rehearsal data for addressing continual VQA. Our proposal, named as GaB, generates pseudo-rehearsal data by posing previous task questions on new task data. Yet, despite being effective, the distribution of generated questions skews towards the most frequently posed questions due to the limited and task-specific training data. To mitigate this issue, we introduce a pseudo-rehearsal balancing module that aligns the generated data towards the ground-truth data distribution using either the question meta-statistics or an unsupervised clustering method. We evaluate our proposed method on two recent benchmarks, i.e. VQACL-VQAv2 and CLOVE-function benchmarks. GaB outperforms all the data-free baselines with substantial improvement in maintaining VQA performance across evolving tasks, while being on-par with methods with access to the past data. Code and models are available at https://github.com/Deepayan137/GaB.
@@ 1 Introduction
We evaluate GaB on continual learning VQA benchmarks, namely VQACL-VQAv2 [57] and CLOVE-function [22] and compare the proposed approach with respect to state-of-the-art and continual learning baselines. Results show a large margin improvement on all considered settings among data-free methods.
@@ 4.1 Comparison
GaB-classifier (Ours) 47.65 3.61
GaB-clustering (Ours) 48.40 1.40
@@ 4.2 Ablation studies
GaB generates questions freely, based on the image content. As we discussed in Sec. 3.2, this results in a skewed distribution towards more prevalent question types and we need to fix this issue via a post-generation balancing strategy. A different direction is to avoid this problem altogether by conditioning the question generation process on target question types. We study this alternative strategy, and compare the proposed approach with GaB-conditioning where the language model is trained by explicitly prompting to generate questions based on their type, p=”Question type <type>: Question: <question> Answer: <answer>”. During inference, question generation is similarly driven by the classifier partitioning function to ensure a balanced distribution. Table 3 presents the results of this conditioning approach in comparison to our methods involving only post-generation processing. We observe that the results of GaB-conditioning are inferior to ours, going from 47.65 to 44.07 in terms of AP. Upon further inspection, we find that forcing the model to produce questions related to certain types can lead to false assumptions, such as forcing to ask a question type that is not relevant for the given image (see Fig. 5 for illustration).
@@
In our method, rehearsal generation simultaneously produces question-answer pairs based on image inputs. Alternatively, we explored using our trained answer generation model to respond to these questions separately in a self-training manner (GaB-self). As detailed in Table 3, this method resulted in a significant decrease in AP by approximately 2.65, likely due to the inferior quality of the pseudo-answers, which are often noisy or incomplete. These issues may cause confirmation bias during fine-tuning. For illustrative examples, see the bottom row of Fig. 5. .
@@
In this study, we evaluate the effectiveness of generating question-answer (QA) pairs with images from the current versus previous tasks. Our goal is to discern why our rehearsal strategy, which uses current task images for QA generation, surpasses traditional methods utilizing real-data. In Table 3, we demonstrate with GaB-pastimages that applying our QA generation strategy to past task images results in a performance decrease of about 10% compared to using current task images, as shown in GaB-classifier.
We believe that using current task images for QA pair generation offers enhanced contextual relevance, similar to data augmentation. This approach not only provides more context by posing multiple questions on the same images but also helps the model bridge knowledge gaps across tasks, effectively serving as a form of positive regularization. Consequently, the model better assimilates new information with previously acquired knowledge, improving its ability to maintain relevant capabilities across different tasks.
@@ 5 Conclusion
6 Acknowledgment
The authors acknowledge the CINECA award under the ISCRA initiative for the availability of high-performance computing resources and support. Deepayan Das is supported by the PRIN project “B-FAIR” (Prot.2022EXF3HX) and the PAT project “AI@TN”. This work was supported by the projects EU Horizon ELIAS (No. 101120237), AI4TRUST (No.101070190), FAIR Future AI Research (PE00000013), funded by NextGeneration EU, and carried out in the Vision and Learning joint laboratory of Fondazione Bruno Kessler and the University of Trento, Italy